# Post-hoc Reward Calibration: A Case Study on Length Bias

### v1

@@ 4.1 Length Calibrated Rewards on RewardBench
We first validate our proposed calibration algorithms using the RewardBench dataset. Each data point in RewardBench comprises one prompt $q$ and two LLMs-generated responses $y_{w}$ and $y_{l}$. The final score is calculated as the averaged accuracy of reward models in identifying $y_{w}$ over four subsets: Chat, Chat Hard, Safety, and Reasoning. We mainly investigate 33 BT-based RMs whose rewarding scores are officially available [^1] https://huggingface.co/datasets/allenai/reward-bench-results and whose accuracy is above 50% (random guess) from the RewardBench Leaderboard as of Aug. 2024.
@@ Results
The overall results for different calibration methods are illustrated in Fig. 1. The line chart on the left showcases the calibrated performance (y-axis) w.r.t original performance (x-axis) of different calibration algorithms for different RMs. The scatter plot on the right highlights the performance gains for different calibration methods. Our empirical findings include: (1) The simple coarse-grained Length Penalty method already leads to a notable improvement (orange line v. blue line), to some extent indicating the prevalence and significance of the RM’s length bias. (2) The reward calibration techniques we introduce yield improvements for nearly all RMs. Among these techniques, the RC-LWR-Penalty achieves a 3.11 performance gain averaged over all 33 models. Though it could lead to performance degradation (5 models for RC-LWR and 4 models for RC-LWR-Penalty), the degradation is very slight and the worst case is only about -1 point for RC-LWR and RC-LWR-Penalty. (3) More importantly, besides improving reward models with moderate performance, the proposed calibration method could also boost the performance of strong reward models. For example, after RC-LWR-Penalty calibration, nine reward models surpass the accuracy of GPT4-Turbo, and the best calibrated model even approaches the latest state-of-the-art (SOTA) LLM-as-Judge models. We also calibrate DPO-based reward models with our proposed method and report results in the Appendix C.
@@ 4.2 Length Calibrated Reward as LLMs Evaluators
A critical application for RM is to serve as a judge to replace human evaluation for evaluating LLMs. Most existing methods (Dubois et al., 2023b; 2024; Li et al., 2024; Kwan et al., 2024) use strong generative LLMs as the judge model. For instance, AlpacaEval (AE1) uses the GPT4’s output logits to calculate the preference probability of the evaluated LLM’s output, $y$, against the responses of the baseline LLM (GPT4-1106-preview), $y_{b}$, for each prompt $q$. This preference probability is then averaged to determine the win rate to rank different LLMs. AlpacaEval2 (AE2) further introduced a length-controlled win rate, a regression-based method to mitigate GPT4’s length bias. In our setting, we primarily use BT-based RMs to score the LLM’s output. Given a prompt $q$ and corresponding output $y$ from the evaluated LLM and output $y_{b}$ from the baseline LLM, the RM scores them as $r$ and $r_{b}$, respectively; then the Bradley-Terry model yields the preference probability $p(y\succ y_{b}|q)=\sigma(r-r_{b})$. We then compute the average of $p(y\succ y_{b}|q)$ across all prompts to determine the final win rate for ranking all evaluated LLMs. We directly use the model’s output from the AlpacaEval official repo. [^2] https://github.com/tatsu-lab/alpaca_eval/tree/main/results Additionally, we apply the length-controlled regression method for our BT-based RMs as one of our baselines, denoted as LC. To further validate the generality of our proposed method, we also apply RC-LWR to AE1. Since AE1 uses GPT4 to compare two responses pairwise and uses its output logits to compute the preference probability $p(y\succ y_{b}|x)$, following Eq. 1, we employ the inverse function of $\operatorname{sigmoid}$ to calculate the reward margin $r-r_{b}$. We then utilise the LWR to directly predict the biased reward margin given the length margin $|y|-|y_{b}|$.
@@
We look at the following metrics following Dubois et al. (2024): (1) Gameability: this measures whether the win rate given by the evaluator is significantly affected by simply promoting the same model to be more concise and verbose. Specifically, five models are selected to be prompted three times: (a) normally, (b) to “give as much detail as possible”, (c) to “be as concise as possible while still providing all the necessary information to answer the question” and get three win rates. Then, the gameability is defined as the normalised variance between those three win rates and averaged across these five models. Higher gameability indicates that the evaluator is more sensitive to the output’s length. (2) Spearman correlation of reward model produced ranking with official AlpacaEval2 leaderboard over 184 models. (3) Spearman correlation of reward model produced ranking with ChatbotArena. We employ three versions of ChatbotArena (CA): CA Feb. is used by Dubois et al. (2024), we introduce it to ensure that we reproduce their results, CA Aug. is the latest one that is more recent and includes more models, and the CA. Length is the CA Aug. disentangled with length proposed by Chiang et al. (2024).
@@
Figure 3: The heatmap demonstrates the Spearman correlation ($\uparrow$) between reward-models-produced rankings with the AlpacaEval2 (AE2) and ChatbotArena (CA). We also report the calibration results for GPT4-1106 on the right, with the Original Reward indicating AlpacaEval1, LC representing AE2, and RC-LWR referring to rankings calibrated with our proposed RC-LWR method. The x-axis represents different calibration methods. The two highest correlations in each row are marked with a dashed rectangle . We select three versions of CA. CA Feb. is used in the paper of AE2 and is selected to validate our reproduction of relevant results. CA Aug. is the latest ranking and contains more models. CA Length is the CA Aug. disentangled with the response length.
@@ Results
The Spearman correlations of RM-produced rankings with the AE2 and the ChatbotArenas are presented in the Fig. 3. Our findings include: (1) Overall, RC-Mean and RC-LWR provide effective calibrations, leading to a higher correlation with AE2 and ChatbotArena, achieving comparable calibration effects with LC on all eight RMs. Note that LC is specifically designed for the AE leaderboard and can not be straightforwardly employed for reward calibration because it requires the model ID and instruction ID for regression, which are not available for general reward calibration. (2) Compared with AE2, RC-LWR calibrated AE achieves comparable Spearman correlations with Chatbot Arena, suggesting that our method can generalise to LLM-as-Judge results. (3) The calibrated GRM-llama3-8B-distill achieves 0.975 with ChatbotArena Feb., and the calibrated FsfairX-LlaMA3-RM-v0.1 achieves 0.975 with AE2 and 0.951 with ChatbotArena Aug., demonstrating the potential of calibrated open-source RM to provide reliable LLM evaluations. (4) The Length Penalty brings improvements in most cases. But it can also lead to degeneration (internlm2-7b-reward) or can be not effective at all (Eurus-RM-7b). This is because different RMs have different reward scales, making it harder to employ the same penalty weight for different RMs. On the contrary, our proposed RC-LWR is more robust to different hyper-parameter choices, rendering it easier to use in practice. (5) Unlike the results on RewardBench, the Length Penalty and RC-LWR-Penalty are less effective in this setting, likely due to differences in length margin ( denoted as $||x_{1}|-|x_{2}||$) distributions of two dataset. For instance, in the RewardBench, 60% of data points have a length margin below 600, while in the AlpacaEval setting, this figure is less than 40%. Consequently, the penalty term has a more significant impact on the AlpacaEval calibration. More relevant analyses are shown in the Appendix C. This underscores the practicality of our method over the penalty-based approach, as the latter may require extensive hyperparameter tuning for penalty weight. Such tuning is influenced not only by the reward model’s scale but also by the output’s length distribution. (6) The rankings calibrated by LC, RC-Mean, RC-LWR are usually more aligned with CA Length compared with CA Aug., further validating the effectiveness of our proposed methods. (7) As shown in Tab. 1, RC-Mean and RC-LWR effectively reduce the Gameability for various reward RMs, demonstrating the calibrated rewards are less sensitive to different prompting techniques and thus more reflective of the response quality.
4.3 Length Calibrated Rewards for LLMs alignment
@@
In this setting, we investigate whether the calibrated rewards could serve as a better training signal for LLMs alignment. Because the post-hoc reward calibration assumes a scored dataset used for calibration, which could be seamlessly integrated with the offline RL, we choose to first employ the RMs to score the on policy dataset and then perform DPO. Our experiment pipeline could be described as follows: for each prompt $q$ in a prompt dataset, multiple responses $y_{i}$ are sampled from an LLM, then scored as $r_{i}$ by an RM. Among these responses, the response with the highest or lowest score is selected to create a preference pair $(p,y_{w},y_{l})$ for Direct Preference Optimisation (DPO) tuning. Specifically, to validate the effectiveness of our method, we choose two instruct models: Meta-Llama-3-8B-Instruct (AI@Meta, 2024) and gemma2-9b-it (Team, 2024) and two reward models: FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2023) and GRM-llama3-8B-sftreg (Yang et al., 2024), resulting in 4 LLM-RM configurations: (1) Llama-3-8B-Fsfairx; (2) Llama-3-8B-GRM; (3) gemma2-9b-Fsfairx; and (4) gemma2-9b-GRM. For each LLM–RM combination, we compare the Original Reward, Length Penalty, and RC-LWR. We directly utilise the on-policy generations333huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm and
@@
of two models generated by Meng et al. (2024). They employ the 60k instructions from the UltraFeedback (Cui et al., 2024) dataset and sample 5 responses for each instruction from the model. Then, we use those two reward models to score the dataset and to conduct DPO tuning. All hyper-parameters related to DPO tuning are set following Meng et al. (2024).
@@
We evaluate the model with the Length-Controlled win rate from the AE2 for open-ended generation. We also test the model on eight popular benchmarks: MMLU (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021), ARC-challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), IFEval (Zhou et al., 2023), GPQA (Rein et al., 2023), and BBH (Suzgun et al., 2023) to see how the proposed calibration affects the benchmark performance. We employ the OpenCompass (Contributors, 2023) for benchmark evaluation and report the average score of these benchmarks. All evaluation configurations are set as a default in OpenCompass.
@@ Results
Results of employing calibrated rewards for LLMs alignment are demonstrated in Tab. 2, where the left and the middle focus on the LC win rate and the right on benchmark performance. We observe that: (1) Similar to the experimental results with the RewardBench, the Length Penalty serves as a simple yet effective baseline, highlighting the significant issue of length exploitation in LLM alignment (Meng et al., 2024). (2) Without extra training and data annotation, the RC-LWR outperforms the Original Reward baseline on AE2 by about 9.5 for Meta-Llama-3-8B-Instruct-based DPO and about 7 points for gemma2-9b-it. It effectively imrpove the output quality while preserving the length. (3) The DPO tuning leads to performance drops on our eight benchmarks. The drop is more significant for gemma2-9b-it-based configurations (e.g., for gemma2-9b-fsfairx, the performance drops from 63.69 to 55.28). This may be because a larger learning rate of 5e-7 is set for gemma2-based experiments compared to 3e-7 for Llama3-based experiments following Meng et al. (2024). However, with the same hyper-parameters, both Length Penalty and RC-LWR methods could mitigate the performance drop, indicating that the length-regularised alignment not only controls the output length but also reserves the model’s benchmark performance during DPO alignment.
Table 2: Results of using length-calibrated rewards for LLMs alignment. Calibration methods we compare include Original Reward (OR), Length Penalty (LP), and RC LWR. We apply these methods in four different LLM–RM configurations. We report the Length-Controlled win rate, the average character length on the AlpacaEval, and average performance on eight benchmarks. The Table illustrates that RC-LWR calibration achieves up to 10% LC win rate improvement and alleviates the performance drop on benchmarks.
@@ 5 Analysis
Does the calibration mitigate the length preference of models?
One desideratum for an ideal calibration method is that the rewards should be weakly correlated with the bias characteristics of interest, thus cancelling the model’s preference for specific characteristics and preventing reward hacking in its downstream applications. Therefore, we calculate the average and variance of the absolute Spearman correlation score over RMs for different methods on the RewardBench dataset and show them in Tab. 3. Our observation includes: (1) Length bias is significant for the board spectrum of RMs (0.293 for original reward). (2) The Length Penalty cannot effectively mitigate the model’s preference for length. The added penalty term renders RM to prefer shorter outputs. (3) In contrast to Length Penalty method, our proposed methods could effectively calibrate the RM’s preference on output length, i.e., the correlation score between output length and reward is approximately 0 for all RMs after calibration.
Calibration behaviours on models with different length preferences.
If an RM exhibits a weak preference for the characteristic of interest, an optimal calibration should not significantly alter its rewarding behaviour and vice versa. Rewarding accuracy alone may not effectively capture this aspect, as a calibration might result in an equal number of correct and incorrect predictions, thereby leaving the overall accuracy unchanged. We thus count how many preference pairs are reversed by the calibration on the RewardBench dataset. We report the fraction of it over the entire dataset. The results are shown in the Fig. 4. Our observations are as follows: (1) Length Penalty demonstrates the moderate correlation between RM length bias and the preference overturned and is unstable. It does not calibrate any pairs for some models with strong positive length preferences. For models with a strong negative preference for length, it is ineffective as the model already prefers shorter outputs in general. (2) Both RC-Mean and RC-LWR methods exhibit a strong correlation: the weaker the bias of the model, the fewer pairs are reversed by the calibration method, indicating that our calibration method will not severely affect the non-biased model. The desirable feature may enhance the practicability of the proposed method.
Figure 4: The number of preferences reversed by the calibration (y-axis) for models with different length preferences (x-axis), measured by the Spearman correlation with length. To determine the correlation between x and y, we further report the Spearman Correlation between them. For RC-Mean and RC-LWR, as the plot is approximately symmetric, we report the correlation between absolute values of x and y.
@@
Figure 5: The calibrated w.r.t original performance of 33 markdown-feature-calibrated BT-based models.
To validate that our proposed method could generalise to other bias features, we apply it to calibrating the bias towards the markdown listing features in the output, motivated by the empirical findings of Dubois et al. (2023a). Observing that LLMs often utilise markdown style lists with bold subheadings as the output format, we define the characteristic function $c(x)$ as the number of headers, lists, and bold strings following  Chiang et al. (2024). The calibration results for 33 BT-based RMs are shown in Fig. 5. We find that calibrating the number of lists in the output also yields improvement on the RewardBench, with a 1.81 averaged performance gain for RC-Mean and 1.86 for RC-LWR. Furthermore, similar to length bias calibration, improvements are observed across nearly all models, and the originally good reward models get slightly stronger after calibration. The gains are not as big as the gains for calibrating length bias, which demonstrates, to some extent, that the length bias is more prevalent.
@@
In Appendix C, we show that the proposed method is robust to different bandwidth choices of LWR (Fig. 7). And its calibration effect can be smoothly adjusted by introducing a constant before the LWR prediction (Fig. 8 and Tab. 4). We also demonstrate that RMs with more substantial bias benefit more from our approach (Fig. 10). We further showcase the potential of the proposed method to calibrate bias from multiple characteristics simultaneously (Fig. 11). Moreover, a stress test evaluating the data efficiency of the method reveals that our method is not significantly constrained by the size of the datasets used for the calibration (Fig. 13).

### v2

@@ 4.1 Length Calibrated Rewards on RewardBench
Each data point in RewardBench comprises one prompt $q$ and two LLMs-generated responses $y_{w}$ and $y_{l}$. The final score is calculated as the averaged accuracy of reward models in identifying $y_{w}$ over four subsets: Chat, Chat Hard, Safety, and Reasoning. We mainly investigate 33 BT-based RMs whose rewarding scores are officially available [^1] https://huggingface.co/datasets/allenai/reward-bench-results and whose accuracy is above 50% (random guess) from the RewardBench Leaderboard as of Aug. 2024.
@@ Results
The overall results for different calibration methods are illustrated in Fig. 1. The line chart on the left showcases the calibrated performance (y-axis) w.r.t original performance (x-axis) of different calibration algorithms for different RMs. The scatter plot on the right highlights the performance gains for different calibration methods. Our empirical findings include: (1) The simple coarse-grained Length Penalty method already leads to a notable improvement (orange line v. blue line), to some extent indicating the prevalence and significance of the RM’s length bias. (2) The reward calibration techniques we introduce yield improvements for nearly all RMs. Among these techniques, the RC-LWR-Penalty achieves a 3.11 performance gain averaged over all 33 models. Though it could lead to performance degradation (5 models for RC-LWR and 4 models for RC-LWR-Penalty), the degradation is very slight and the worst case is only about -1 point for RC-LWR and RC-LWR-Penalty. (3) More importantly, besides improving reward models with moderate performance, the proposed calibration method could also boost the performance of strong reward models. For example, after RC-LWR-Penalty calibration, nine reward models surpass the accuracy of GPT4-Turbo, and the best calibrated model even approaches the latest state-of-the-art (SOTA) LLM-as-Judge models. We also calibrate DPO-based reward models with our proposed method and report results in the Appendix D.
@@ 4.2 Length Calibrated Reward as LLMs Evaluators
The typical LLM evaluation pipeline first constructs a prompt dataset, samples responses from different LLMs, and then calculates the win rate between different models. Based on this pipeline, AlpacaEval (AE1) (Dubois et al., 2023b), with GPT4-1106-Preview as the evaluator, fixes a baseline LLM and uses the win rates against the same baseline LLM to produce the ranking. To calibrate the length bias of GPT4 evaluation, AlpacaEval2 (AE2) (Dubois et al., 2024) introduced length-controlled regression to calibrate the win rate of AE1 and became a popular generative LLM leaderboard. Based on the AE1 prompt dataset, this section tests whether the open-source BT-based RMs could provide a more reliable evaluation for LLMs after calibration. Because the length-controlled technique proposed in AE2 can apply to BT-based RMs, we include it in our baselines, denoted as LC. We also apply our method to calibrate AE1 to further compare with LC. More details about applying RM for AE and calibrating the produced ranking are in Appendix C.
@@
Following Dubois et al. (2024), we look at the following metrics to see whether the calibrated rewards could provide a more reliable evaluation: (1) Gameability: this measures whether the win rate given by the evaluator is significantly affected by simply promoting the same model to be more concise or verbose. Like Dubois et al. (2024), five LLMs are selected to be prompted to generate the response normally, verbosely, and concisely. Then, the gameability is defined as the normalised variance between those three win rates and averaged across these five models. Higher gameability indicates that the evaluator is more sensitive to the output’s length and could help us identify evaluators with length bias. (2) Spearman correlation with existing popular LLM leaderboard, including AE2 and ChatbotArena (CA). We employ three versions of (CA): (i) CA Feb., used by Dubois et al. (2024), to reproduce their results; (ii) CA Aug., the latest version that includes more LLMs; and (iii) CA. Length, namely CA Aug. calibrated with respect to length as proposed by Chiang et al. (2024).
@@
Figure 3: The heatmap demonstrates the Spearman correlation ($\uparrow$) between reward-models-produced rankings with the AlpacaEval2 (AE2) and ChatbotArena (CA). We also report the calibration results for GPT4-1106 on the right, with the Original Reward indicating AlpacaEval1, LC representing AE2, and RC-LWR referring to rankings calibrated with our proposed RC-LWR method. The x-axis represents different calibration methods. The two highest correlations in each row are marked with a dashed rectangle . We select three versions of CA: (i) CA Feb., from the AE2 paper, to validate our reproduction of the original results; (ii) CA Aug., the latest ranking which contains more LLMs; and (iii) CA Length, which is CA Aug. disentangled with respect to the response length.
@@ Results
The Spearman correlations of RM-produced rankings with the AE2 and the ChatbotArenas are presented in the Fig. 3. Our findings include: (1) Overall, RC-Mean and RC-LWR provide effective calibrations, leading to a higher correlation with AE2 and ChatbotArena, achieving comparable calibration effects with LC on all eight RMs. Note that LC is specifically designed for the AE leaderboard and can not be straightforwardly employed for reward calibration because it requires the LLM ID and instruction ID for regression, which are not available for general reward calibration. (2) Compared with AE2, RC-LWR calibrated AE achieves comparable Spearman correlations with Chatbot Arena, suggesting that our method can generalise to LLM-as-Judge results. (3) The calibrated GRM-llama3-8B-distill achieves 0.975 with ChatbotArena Feb., and the calibrated FsfairX-LlaMA3-RM-v0.1 achieves 0.975 with AE2 and 0.951 with ChatbotArena Aug., demonstrating the potential of calibrated open-source RM to provide reliable LLM evaluations. (4) The Length Penalty brings improvements in most cases. But it can also lead to degeneration (internlm2-7b-reward) or can be not effective at all (Eurus-RM-7b). This is because different RMs have different reward scales, making it harder to employ the same penalty weight for different RMs. On the contrary, our proposed RC-LWR is more robust to different hyper-parameter choices, rendering it easier to use in practice. (5) Unlike the results on RewardBench, the Length Penalty and RC-LWR-Penalty are less effective in this setting, likely due to differences in length margin ( denoted as $||x_{1}|-|x_{2}||$) distributions of two dataset. For instance, in the RewardBench, 60% of data points have a length margin below 600, while in the AlpacaEval setting, this figure is less than 40%. Consequently, the penalty term has a more significant impact on the AlpacaEval calibration. More relevant analyses are shown in the Appendix D. This underscores the practicality of our method over the penalty-based approach, as the latter may require extensive hyperparameter tuning for penalty weight. Such tuning is influenced not only by the RM’s reward scale but also by the output’s length distribution. (6) The rankings calibrated by LC, RC-Mean, RC-LWR are usually more aligned with CA Length compared with CA Aug., further validating the effectiveness of our proposed methods. (7) As shown in Tab. 1, RC-Mean and RC-LWR effectively reduce the Gameability for various reward RMs, demonstrating the calibrated rewards are less sensitive to different prompting techniques and thus more reflective of the response quality.
4.3 Length Calibrated Rewards for LLMs’ alignment
@@
In this setting, we investigate whether the calibrated rewards could serve as a better training signal for LLMs’ alignment. Because the post-hoc reward calibration assumes a scored dataset used for calibration, which could be seamlessly integrated with the offline RL, we choose to first employ the RMs to score the on policy dataset and then perform DPO. Our experiment pipeline could be described as follows: for each prompt $q$ in a prompt dataset, multiple responses $y_{i}$ are sampled from an LLM, then scored as $r_{i}$ by an RM. Among these responses, the response with the highest or lowest score is selected to create a preference pair $(p,y_{w},y_{l})$ for Direct Preference Optimisation (DPO) tuning. Specifically, to validate the effectiveness of our method, we choose two instruct LLMs: Meta-Llama-3-8B-Instruct (AI@Meta, 2024) and gemma2-9b-it (Team, 2024) and two reward models: FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2023) and GRM-llama3-8B-sftreg (Yang et al., 2024), resulting in 4 LLM-RM configurations: (1) Llama-3-8B-Fsfairx; (2) Llama-3-8B-GRM; (3) gemma2-9b-Fsfairx; and (4) gemma2-9b-GRM. For each LLM–RM combination, we compare the Original Reward, Length Penalty, and RC-LWR. We directly utilise the on-policy generations222huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm and
@@
of two selected LLMs generated by Meng et al. (2024). They employ the 60k instructions from the UltraFeedback (Cui et al., 2024) dataset and sample 5 responses for each instruction from the LLM. Then, we use those two RMs to score the dataset and to conduct DPO tuning. All hyper-parameters related to DPO tuning are set following Meng et al. (2024).
@@
We evaluate the LLM with the Length-Controlled win rate from the AE2 for open-ended generation. We also test the LLM on eight popular benchmarks: MMLU (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021), ARC-challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), IFEval (Zhou et al., 2023), GPQA (Rein et al., 2023), and BBH (Suzgun et al., 2023) to see how the proposed calibration affects the benchmark performance. We employ the OpenCompass (Contributors, 2023) for benchmark evaluation and report the average score of these benchmarks. All evaluation configurations are set as a default in OpenCompass.
@@ Results
Results of employing calibrated rewards for LLMs’ alignment are demonstrated in Tab. 2, where the left and the middle focus on the LC win rate and the right on benchmark performance. We observe that: (1) Similar to the experimental results with the RewardBench, the Length Penalty serves as a simple yet effective baseline, highlighting the significant issue of length exploitation in LLM alignment (Meng et al., 2024). (2) Without extra training and data annotation, the RC-LWR outperforms the Original Reward baseline on AE2 by about 9.5 for Meta-Llama-3-8B-Instruct-based DPO and about 7 points for gemma2-9b-it. It effectively improves the output quality while preserving the length. (3) The DPO tuning leads to performance drops on our eight benchmarks. The drop is more significant for gemma2-9b-it-based configurations (e.g., for gemma2-9b-fsfairx, the performance drops from 63.69 to 55.28). This may be because a larger learning rate of 5e-7 is set for gemma2-based experiments compared to 3e-7 for Llama3-based experiments following Meng et al. (2024). However, with the same hyper-parameters, both Length Penalty and RC-LWR methods could mitigate the performance drop, indicating that the length-regularised alignment not only controls the output length but also reserves the LLM’s benchmark performance during DPO alignment.
Table 2: Results of using length-calibrated rewards for LLMs’ alignment. Calibration methods we compare include Original Reward (OR), Length Penalty (LP), and RC LWR. We apply these methods in four different LLM–RM configurations. We report the Length-Controlled win rate, the average character length on the AlpacaEval, and average performance on eight benchmarks. The Table illustrates that RC-LWR calibration achieves up to 10% LC win rate improvement and alleviates the performance drop on benchmarks.
@@ 5 Analysis
Does the calibration mitigate the length preference of RMs?
One desideratum for an ideal calibration method is that the rewards should be weakly correlated with the bias characteristics of interest, thus cancelling the RM’s preference for specific characteristics and preventing reward hacking in its downstream applications. Therefore, we calculate the average and variance of the absolute Spearman correlation score over RMs for different methods on the RewardBench dataset and show them in Tab. 3. Our observation includes: (1) Length bias is significant for the broad spectrum of RMs (0.293 for original reward). (2) The Length Penalty cannot effectively mitigate the RM’s preference for length. The added penalty term renders RM to prefer shorter outputs. (3) In contrast to Length Penalty method, our proposed methods could effectively calibrate the RM’s preference on output length, i.e., the correlation score between output length and reward is approximately 0 for all RMs after calibration.
Calibration behaviours on RMs with different length preferences.
If an RM exhibits a weak preference for the characteristic of interest, an optimal calibration should not significantly alter its rewarding behaviour and vice versa. Rewarding accuracy alone may not effectively capture this aspect, as a calibration might result in an equal number of correct and incorrect predictions, thereby leaving the overall accuracy unchanged. We thus count how many preference pairs are reversed by the calibration on the RewardBench dataset. We report the fraction of it over the entire dataset. The results are shown in the Fig. 4. Our observations are as follows: (1) Length Penalty demonstrates the moderate correlation between RM length bias and the preference overturned and is unstable. It does not calibrate any pairs for some RMs with strong positive length preferences. For RMs with a strong negative preference for length, it is ineffective as the RM already prefers shorter outputs in general. (2) Both RC-Mean and RC-LWR methods exhibit a strong correlation: the weaker the bias of the RM, the fewer pairs are reversed by the calibration method, indicating that our calibration method will not severely affect the non-biased RM. The desirable feature may enhance the practicability of the proposed method.
Figure 4: The number of preferences reversed by the calibration (y-axis) for RMs with different length preferences (x-axis), measured by the Spearman correlation with length. To determine the correlation between x and y, we further report the Spearman Correlation between them. For RC-Mean and RC-LWR, as the plot is approximately symmetric, we report the correlation between absolute values of x and y.
@@
Figure 5: RewardBench score before (x-axis) and after (y-axis) markdown-feature-calibration.
To validate that our proposed method could generalise to other bias, we apply it to calibrating the bias towards the markdown listing features in the output, motivated by the empirical findings of Dubois et al. (2023a). Observing that LLMs often utilise markdown style lists with bold subheadings as the output format, we define the characteristic function $c(x)$ as the number of headers, lists, and bold strings following  Chiang et al. (2024). The calibration results for 33 BT-based RMs are shown in Fig. 5. We find that calibrating the number of lists in the output also yields improvement on the RewardBench, with a 1.81 averaged performance gain for RC-Mean and 1.86 for RC-LWR. Furthermore, similar to length bias calibration, improvements are observed across nearly all RMs, and the originally good RMs get slightly stronger after calibration. The gains are not as big as the gains for calibrating length bias, which demonstrates, to some extent, that the length bias is more prevalent.
@@
In Appendix D, we show that the proposed method is robust to different hyper-parameter choices of LWR (Fig. 7). For the cases when the independence assumption does not hold, we introduce a calibration constant before LWR prediction to determine how much bias to subtract, thus smoothly controlling the calibration effect (Fig. 8 and Tab. 4). We also demonstrate that RMs with more substantial bias benefit more from our approach (Fig. 10). We further showcase the potential of the proposed method to calibrate bias from multiple characteristics simultaneously (Fig. 11). Moreover, a stress test evaluating the data efficiency of the method reveals that our method is not significantly constrained by the size of the datasets used for the calibration (Fig. 13).